# Robust Robotic Exploration and Mapping Using Generative Occupancy Map Synthesis

### v1

@@ Abstract
We present a novel approach for enhancing robotic exploration by using generative occupancy mapping. We introduce SceneSense, a diffusion model designed and trained for predicting 3D occupancy maps given partial observations. Our proposed approach probabilistically fuses these predictions into a running occupancy map in real-time, resulting in significant improvements in map quality and traversability. We implement SceneSense onboard a quadruped robot and validate its performance with real-world experiments to demonstrate the effectiveness of the model. In these experiments we show that occupancy maps enhanced with SceneSense predictions better represent our fully observed ground truth data ($24.44\%$ FID improvement around the robot and $75.59\%$ improvement at range). We additionally show that integrating SceneSense-enhanced maps into our robotic exploration stack as a “drop-in” map improvement, utilizing an existing off-the-shelf planner, results in improvements in robustness and traversability time. Finally we show results of full exploration evaluations with our proposed system in two dissimilar environments and find that locally enhanced maps provide more consistent exploration results than maps constructed only from direct sensor measurements.
@@ 1 Introduction
Efficient and reliable robotic exploration is fundamental to the practical deployment of robotic systems for use in real-world scenarios. In order to intelligently explore, an autonomous agent must have a useful representation of its surroundings and be able to localize within the environment. Robotic agents deployed in search and rescue [1, 2] or disaster response missions [3, 4, 5] rarely have the environment mapped a priori. The robotic agent must operate cautiously as it builds the map from direct sensor observations. This results in slower, less efficient exploration as the agent can only operate over terrain in the sensor field-of-view and large portions of the space may be obscured from that view.
@@
This lack of prior scene information means agents are constantly operating on incomplete information. In addition to limited field-of-view for mapping, there is also sensor noise, unpredictable dynamic obstacles, obstruction of key information, and visual degradation that an agent may need to reason over to complete a given task. While human agents can filter out noise and fill in gaps in observations by drawing on prior experiences, robots are generally limited to using direct observations without the benefits of prior knowledge or intuition. This results in robotic agents that only take action based on direct observations, a process that further reduces exploration efficiency because of time consuming replanning when informative sensor updates occur.
@@
In related fields, such as self-driving cars, researchers attempt to eliminate the mapping uncertainty using pre-compiled high-definition (HD) maps [6]. This approach provides a good method of localization at the global level, but there still exists uncertainty and missing information locally. Furthermore, this approach scales poorly, as new HD maps are required for any new operating environment. As discussed previously there are many applications where collating an HD map before navigation is not possible, such as in search and rescue scenarios [7] or extra-terrestrial exploration [8].
@@
With recent breakthroughs in generative AI, neural networks are able to generate convincing human-like results in various modalities, such as text [9], images [10, 11], and audio [12]. We leverage these models to enable robots to make common-sense inference of geometry that is occluded from view. Our model enables a robotic agent to infer geometry through occupancy prediction beyond an obstruction so that the agent can plan a trajectory beyond the field-of-view similar to the way a human would navigate occlusions using intuition.
In this work, we use the occupancy prediction diffusion network SceneSense [13, 14] to generate occupancy information in key areas during exploration. We design a robotic system to run SceneSense online during exploration and develop the probabilistic update rules to merge the diffusion information with a running observed map. We show that with key modifications we can use SceneSense to generate realistic occupancy predication at any location in a map, rather than just local to the platform. Furthermore, we show the enhanced maps that merge the predicted occupancy maps simplify robotic exploration, solving key perennial issues for autonomous platforms that previously required tedious fine tuning of parameters.
@@
We generate results on the efficacy of these methods through field tests with a quadruped platform. We compare robot-centric performance to frontier-centric performance, and we provide a comparison of one-shot map merging to probabilistic map merging. In addition, our testing shows successful navigation in several scenarios which previously challenged robotic perception systems including occupancy mapping at startup, narrow hallways, and glass railings. We provide detailed insights into the system specifications, compute stack, training times, and experiment design. Code used for replicating this work is available at [URL-added-post-publication].
@@ 2.1 Scene Exploration
An autonomous agent must have a useful representation of its surroundings and be able to localize within the environment to efficiently and intelligently explore. A proven strategy for mapping complex environments for use with autonomous systems is 3D volumetric mapping [15, 16]. OctoMap [17] is a popular framework for 3D mapping due to its memory efficiency and wide adoption in the robotics community [18]. In the DARPA Subterranean (SubT) Challenge [7], Team MARBLE [19] utilized OctoMap on a Boston Dynamics Spot robot and finished 3rd overall.
@@
Figure 1: The reverse diffusion process takes the local occupancy information and the Gaussian noise of the area to be diffused over. Noise commensurate with the current diffusion step is added to the local occupancy information, which includes occupied (green) and observed unoccupied (red) data. The result is inpainted into the noisy local occupancy prediction. The inpainted noise data is provided to the denoising network which generates a new noisy geometry prediction at $t-1$. This processes is repeated as the starting noise $x_{T}$ is iteratively denoised to $x_{0}$ which is the final geometry prediction from the framework. This process is further detailed in Algorithm 1.
@@
Other works [20, 21, 22, 23] have shown considerable success using frontier based exploration strategies. In [24], a frontier-based approach is utilized to demonstrate the efficacy of biasing towards unseen areas of a map for more effective mobile robot exploration. Further efficiency improvements to the frontier based exploration strategy are introduced in [25]. In [23], a frontier based method is proposed for safe and efficient map-building by defining a “safe region” around the robot, allowing it to explore and gather new information while ensuring obstacle-free navigation, although it is limited to mapping at a fixed height, making it unsuitable for more complex environments.
@@
In [26], a volumetric map is incrementally generated from depth sensors and a graph based planner [18] reasons over that map to determine if the terrain is traversable for each vertex of a four-legged robot. This graph traversal requires considerable computation time for determining the status of voxels beneath and around the robot platform. Additionally, this method does not address the issues with path planning if there are obstacles or geometric occlusions in the sensor field of view.
@@
Despite the improvements in mapping and frontier based exploration strategies, Team MARBLE [19] reported that there are still challenges with efficient exploration. During the SubT challenge, exploring robots frequently wasted valuable time attempting to explore in previously seen areas rather than extending their trajectories beyond their current frontiers [19]. Influxes of new information from onboard sensors that came when turning corners or exploring new rooms caused the robots to pause for extended periods of time for re-planning, slowing the pace of exploration considerably [19]. The challenge outcomes demonstrated the utility of 3D volumetric mapping using frontier-based exploration for discovery, but also elucidated the need for increased efficiency in the planning and exploration process for time critical missions. Our novel approach improves the efficiency of exploration by reducing ambiguity that leads to long re-planning delays for the robots.
@@ 2.2 Occupancy Prediction
Wang et al. [27] attempts to address challenges of autonomous navigation in environments with occlusions by utilizing a deep learning-based approach to predict the occupancy distribution of portions of the environment. Their method, the Occupancy Prediction Network (OPNet), predicts obstacle distributions, utilizing a self-supervised learning technique that generates training data through simulated navigation trajectories with simulated sensor noise. This approach relies on removing portions of the data from the Matterport3D [28] dataset to train the neural network which biases the model to predict occupancy only for the types of occlusions removed from the original dataset. This approach does not scale well to large unseen sections of an environment, is limited to static updates, and does not generalize due to its reliance on a specially curated dataset.
More recently, Huang et al. [29] introduced SelfOcc, a self-supervised method for 3D occupancy prediction using video sequences. SelfOcc converts 2D images into 3D representations with deformable attention layers and uses signed distance fields (SDF) for regularization and occupancy boundary determination. Although SelfOcc performs well with surrounding cameras, it cannot predict occupancy beyond the camera view due to the limitations of SDF in handling occluded geometry.
@@ 2.3 Scene Synthesis
Diffusion models [30, 31], are a popular and promising technique which have demonstrated impressive results in image [10], video [32], and natural language [33]. Based on these successes, diffusion models are being extended to 3D scene and shape generation. Recent work [34] demonstrates the use of diffusion models for 3D point cloud generation for simple shapes and objects (e.g. tables, chairs). Kim et al. [35] shows successful 3D shape generation from 2D content such as images, and Vahdat et al. [36] demonstrates similar 3D shape generation but using point cloud datasets rather than images. In LegoNet [37], diffusion models are used to rearrange objects in a 3D scene. In DiffuScene [38], a denoising diffusion model is used with text conditioning to generate 3D indoor scenes from sets of unordered object attributes. Unlike these previous works which primarily focus on generating simple shapes, rearranging objects, or creating indoor scenes, our approach leverages diffusion models to fuse generated terrain with the local robot field of view, thereby bridging the gap between 3D scene generation and practical robotics applications.
@@ 2.4 Generative AI in Robotics
The body of work devoted to applications of generative AI in robotics is steadily growing. Yuan et al. [39] demonstrates that hierarchical generative modeling, inspired by human motor control, can enable autonomous robots to effectively perform complex tasks with robust performance even in challenging conditions. DALL-E-Bot [11], uses web-scale diffusion models for generating an image from a text prompt which the robot utilizes to rearrange real objects in accordance with the image. Diffusion models were also used in [40] to improve robot motion planning by learning priors on trajectory distributions from previously successful plans, a methodology which shows good generalization capabilities in environments with previously unseen obstacles. More recently, diffusion models were shown to successfully generate terrain predictions behind occluded geometries in indoor environments using a single RGB sensor mounted on a mobile robot platform [13]. We build on this work, extending the practical applications of diffusion models by demonstrating a novel method for occupancy prediction which shows significant efficiency gains during exploration.
@@ 3 Problem Statement and Preliminaries
The goal of this work is to create and evaluate a method for dense occupancy prediction which can provably enhance a robotic platforms ability to navigate. The ability to generate occupancy predictions beyond occlusions and at range from a mobile platform introduces the subsequent challenge of how to identify and select frontier points for exploration in these generated occupancy spaces. We present an approach to frontier identification and selection for exploration as part of this work.
@@ 3.1 Autonomous Robotic Exploration
Let $\mathbb{M}$ be the current occupancy map, built via measurements from an onboard sensor $\mathbb{S}$ and odometer measurements $\mathbb{O}$. The map consists of voxels $m$ that are categorized as $m\in\mathbb{M}_{free}$, $m\in\mathbb{M}_{occupied}$, or $m\in\mathbb{M}_{unknown}$, representing free, occupied and unknown space respectively. During Robotic Exploration we seek to generate the complete map $\mathbb{M}_{comp}$ of a target area defined as $\mathbb{M}_{target}$ constructed from successive sensor measurements from an onboard sensor $\mathbb{S}$ and odometer measurements $\mathbb{O}$. While completion and alignment of $\mathbb{M}_{comp}$ when compared to $\mathbb{M}_{target}$ is the primary goal of robotic exploration, speed of exploration and naturally by extension volumetric gain per second are secondary goals during autonomous exploration, so long as these goal do not compromise the safety of the system.
@@ 3.2 Frontier Identification and Evaluation
We identify and evaluate frontiers of interest in $\mathbb{M}$ that can enhance the robot’s understanding of the environment during exploration. In general, “interesting” frontiers will maximize the number of unknown voxels available for occupancy prediction while considering common exploration metrics such a directionality, distance from target, and reachability [19, 18].
@@ 3.3 Dense Occupancy Prediction
Dense occupancy prediction predicts the occupancy from $[0,1]$ where 0 is unoccupied and 1 is occupied for every voxel $m$ in a target region $x\subset\mathbb{M}$.
@@ 3.4 Forward Diffusion
$x_{0}$ is defined as a clean occupancy grid where the distribution of $x_{0}$ can be defined as $q(x_{0})$. By sampling from the data distribution $x_{0}\sim q(x_{0})$ the forward diffusion process is defined as a Markov chain of variables $x_{1},...,x_{T}$ that iteratively adds Gaussian noise to the sample. A diffusion step at time $t$ in this chain is defined as:
@@
Conveniently we can apply the reparameterization trick to directly sample $x_{t}$ given $x_{0}$ using the conditional distribution:
@@ 3.5 Reverse Diffusion
Reverse diffusion is a Markov chain of learned Gaussian transitions $p_{\theta}(x_{t-1}|x_{t})$ which is parameterized by a learnable network $\theta$:
@@
where $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the predicted mean and covariance respectively of the Gaussian $x_{t-1}$. Given the initial state of a noisy occupancy map from a standard multivariate Gaussian distribution $x_{t}\sim\mathcal{N}(0,I)$ the reverse diffusion process iteratively predicts $x_{t-1}$ at each time step $t$ until reaching the final state $x_{0}$ which is the goal occupancy map. Similar to the Markov chain defined forward diffusion process the joint distribution of the reverse diffusion process is simply the product of the applied learned Gaussian transitions $p_{\theta}(x_{t-1}|x_{t})$:

### v2

@@ Abstract
We present a novel approach for enhancing robotic exploration by using generative occupancy mapping. We implement SceneSense, a diffusion model designed and trained for predicting 3D occupancy maps given partial observations. Our proposed approach probabilistically fuses these predictions into a running occupancy map in real-time, resulting in significant improvements in map quality and traversability. We deploy SceneSense on a quadruped robot and validate its performance with real-world experiments to demonstrate the effectiveness of the model. In these experiments we show that occupancy maps enhanced with SceneSense predictions better estimate the distribution of our fully observed ground truth data ($24.44\%$ FID improvement around the robot and $75.59\%$ improvement at range). We additionally show that integrating SceneSense enhanced maps into our robotic exploration stack as a “drop-in” map improvement, utilizing an existing off-the-shelf planner, results in improvements in robustness and traversability time. Finally, we show results of full exploration evaluations with our proposed system in two dissimilar environments and find that locally enhanced maps provide more consistent exploration results than maps constructed only from direct sensor measurements.
@@ 1 Introduction
Efficient and reliable robotic exploration is fundamental to the practical deployment of robotic systems for use in real-world scenarios. In order to intelligently explore, an autonomous agent must have a useful representation of its surroundings and be able to localize within the environment. Robotic agents deployed in search and rescue [schwarz2017nimbro, bernard2011autonomous] or disaster response missions [kawatsuma2012emergency, park2017disaster, kruijff2021germanemergencyresponse] rarely have the environment mapped a priori. The robotic agent must operate cautiously as it builds the map from direct sensor observations. This results in slower, less efficient exploration as the agent can only operate over terrain in the sensor field-of-view and large portions of the space may be obscured from that view.
@@
This lack of prior scene information means agents are constantly operating on incomplete information. In addition to limited field-of-view for mapping, there are sensor noise issues, unpredictable dynamic obstacles, obstructions of key information, and visual degradation that an agent may need to reason over to complete a given task. While human agents can filter out noise and fill in gaps in observations by drawing on prior experiences, robots are generally limited to using direct observations without the benefits of prior knowledge or intuition. This results in robotic agents that only take action based on direct observations, a process that further reduces exploration efficiency because of time consuming replanning when informative sensor updates occur.
@@
In related fields, such as self-driving cars, researchers attempt to eliminate the mapping uncertainty using pre-compiled high-definition (HD) maps [ma2019exploiting]. This approach provides a good method of localization at the global level, but there still exists uncertainty and missing information locally. Furthermore, this approach scales poorly, as new HD maps are required for any new operating environment. As discussed previously there are many applications where collating an HD map before navigation is not possible, such as in search and rescue scenarios [chung2023into] or extra-terrestrial exploration [Verma_2024_perseverance_mars].
@@
With recent breakthroughs in generative AI, neural networks are able to generate convincing human-like results in various modalities, such as text [touvron2023llama], images [rombach2022high, kapelyukh2023DALLEBOT], and audio [dhariwal2020jukebox]. We leverage these models to enable robots to make common-sense inference of geometry that is occluded from view. Our model enables a robotic agent to infer geometry through occupancy prediction beyond an obstruction so that the agent can plan a trajectory beyond the field-of-view similar to the way a human would navigate occlusions using intuition.
SceneSense was first introduced in [reed2024scenesense], where a diffusion-based model was used to generate voxelized predictions of occupancy. Subsequent work improved the efficiency of SceneSense, notably achieving faster training and inference times [reed2024online]. These works established that diffusion models can be used for occupancy prediction but left open questions regarding real-time deployment and map level reasoning.
In this work, we extend the SceneSense framework in three ways. First, we demonstrate real-time occupancy generation on a quadruped platform using a diffusion network. Second, we integrate a probabilistic update rule that fuses diffusion based predictions with an evolving observed map, allowing predictions to accumulate and refine over time. Third, we show that SceneSense can generate realistic occupancy predictions across the full map rather than being restricted to regions local to the robot. This broader predictive capability leads to enhanced global maps which can simplify exploration in real world deployments.
@@
We validate our approach through field experiments with a quadruped platform. Specifically, we (i) compare robot-centric and frontier-centric exploration performance, (ii) evaluate one-shot versus probabilistic map merging, and (iii) demonstrate successful navigation in challenging scenarios that historically degrade robotic perception (e.g. transparent structures, narrow hallways). Finally, we provide detailed reporting of system specifications, compute requirements, training times, experimental design, and system limitations. Code for reproducing our results will be made available at [URL-added-post-publication].
@@ 2.1 Scene Exploration
An autonomous agent must have a useful representation of its surroundings and be able to localize within the environment to efficiently and intelligently explore. A proven strategy for mapping complex environments for use with autonomous systems is 3D volumetric mapping [ahmad20213dunstructured, ahmad2021reactiveobstacleavoidance]. OctoMap [hornung13octomap] is a popular framework for 3D mapping due to its memory efficiency and wide adoption in the robotics community [dang2020graph]. In the DARPA Subterranean (SubT) Challenge [chung2023into], Team MARBLE [biggie2023flexible] utilized OctoMap on a Boston Dynamics Spot robot and finished 3rd overall.
@@
Figure 1: The reverse diffusion process takes the local occupancy information and the Gaussian noise of the area to be diffused over. Noise commensurate with the current diffusion step is added to the local occupancy information, which includes occupied (green) and observed unoccupied (red) data. The result is inpainted into the noisy local occupancy prediction. The inpainted noise data is provided to the denoising network which generates a new noisy geometry prediction at t1t-1. This process is repeated as the starting noise $x_{T}$ is iteratively denoised to $x_{0}$ which is the final geometry prediction from the framework. This process is further detailed in Algorithm 1.
@@
Other works [yamauchi1997frontier, yamauchi1998mobile, makarenko2002experiment, gonzalez2002navigation] have shown considerable success using frontier-based exploration strategies. In [freda2005frontier], a frontier-based approach is utilized to demonstrate the efficacy of biasing towards unseen areas of a map for more effective mobile robot exploration. Further efficiency improvements to the frontier based exploration strategy are introduced in [keidar2014efficient]. In [gonzalez2002navigation], a frontier based method is proposed for safe and efficient map-building by defining a “safe region” around the robot, allowing it to explore and gather new information while ensuring obstacle-free navigation, although it is limited to mapping at a fixed height, making it unsuitable for more complex environments.
@@
In [CERBERUStranzatto2022team], a volumetric map is incrementally generated from depth sensors and a graph-based planner [dang2020graph] reasons over that map to determine if the terrain is traversable for each vertex of a four-legged robot. This graph traversal requires considerable computation time for determining the status of voxels beneath and around the robot platform. Additionally, this method does not address the issues with path planning if there are obstacles or geometric occlusions in the sensor field of view.
@@
Despite the improvements in mapping and frontier-based exploration strategies, Team MARBLE [biggie2023flexible] reported that there are still challenges with efficient exploration. During the SubT challenge, exploring robots frequently wasted valuable time attempting to explore in previously seen areas rather than extending their trajectories beyond their current frontiers [biggie2023flexible]. Influxes of new information from onboard sensors that came when turning corners or exploring new rooms caused the robots to pause for extended periods of time for re-planning, slowing the pace of exploration considerably [biggie2023flexible]. The challenge outcomes demonstrated the utility of 3D volumetric mapping using frontier-based exploration for discovery, but also elucidated the need for increased efficiency in the planning and exploration process for time critical missions. Our novel approach improves the efficiency of exploration by reducing ambiguity that leads to long re-planning delays for the robots.
@@ 2.2 Occupancy Prediction
wang2021learning3doccupancyprediction attempts to address challenges of autonomous navigation in environments with occlusions by utilizing a deep learning-based approach to predict the occupancy distribution of portions of the environment. Their method, the Occupancy Prediction Network (OPNet), predicts obstacle distributions, utilizing a self-supervised learning technique that generates training data through simulated navigation trajectories with simulated sensor noise. This approach relies on removing portions of the data from the Matterport3D [chang2017matterport3d] dataset to train the neural network which biases the model to predict occupancy only for the types of occlusions removed from the original dataset. This approach does not scale well to large unseen sections of an environment, is limited to static updates, and does not generalize due to its reliance on a specially curated dataset.
More recently, Huang_2024_CVPR introduced SelfOcc, a self-supervised method for 3D occupancy prediction using video sequences. SelfOcc converts 2D images into 3D representations with deformable attention layers and uses signed distance fields (SDF) for regularization and occupancy boundary determination. Although SelfOcc performs well with surrounding cameras, it cannot predict occupancy beyond the camera view due to the limitations of SDF in handling occluded geometry.
@@ 2.3 Scene Synthesis
Diffusion models [sohl2015deep, ho2020denoising], are a popular and promising technique which have demonstrated impressive results in image [rombach2022high], video [harvey2022flexible], and natural language [huang2022prodiff]. Based on these successes, diffusion models are being extended to 3D scene and shape generation. Recent work [luo2021diffusionpointcloudgeneration] demonstrates the use of diffusion models for 3D point cloud generation for simple shapes and objects (e.g. tables, chairs). kim2023scenegendiffusionmodels shows successful 3D shape generation from 2D content such as images, and vahdat2022lion demonstrates similar 3D shape generation but using point cloud datasets rather than images. In LegoNet [yu2022legonet], diffusion models are used to rearrange objects in a 3D scene. In DiffuScene [tang2023diffuscene], a denoising diffusion model is used with text conditioning to generate 3D indoor scenes from sets of unordered object attributes. Unlike these previous works which primarily focus on generating simple shapes, rearranging objects, or creating indoor scenes, our approach leverages diffusion models to fuse generated terrain with the local robot field of view, thereby bridging the gap between 3D scene generation and practical robotics applications.
@@ 2.4 Generative AI in Robotics
The body of work devoted to applications of generative AI in robotics is steadily growing. yuan2023hierarchicalgenmodeling demonstrates that hierarchical generative modeling, inspired by human motor control, can enable autonomous robots to effectively perform complex tasks with robust performance even in challenging conditions. DALL-E-Bot [kapelyukh2023DALLEBOT], uses web-scale diffusion models for generating an image from a text prompt which the robot utilizes to rearrange real objects in accordance with the image. Diffusion models were also used in [carvalho2023planningdiffusionmodels] to improve robot motion planning by learning priors on trajectory distributions from previously successful plans, a methodology which shows good generalization capabilities in environments with previously unseen obstacles. More recently, diffusion models were shown to successfully generate terrain predictions behind occluded geometries in indoor environments using a single RGB sensor mounted on a mobile robot platform [reed2024scenesense]. We build on this work, extending the practical applications of diffusion models by demonstrating a novel method for occupancy prediction which shows significant efficiency gains during exploration.
@@ 3 Problem Statement and Preliminaries
The goal of this work is to create and evaluate a method for dense occupancy prediction that demonstrably enhances a robotic platform’s ability to navigate, as measured by reducing the time required to fully explore its environment. The ability to generate occupancy predictions beyond occlusions and at range from a mobile platform introduces the subsequent challenge of how to identify and select frontier points for exploration in these generated occupancy spaces. We present an approach to frontier identification and selection for exploration as part of this work.
@@ 3.1 Autonomous Robotic Exploration
Let $\mathbb{M}$ denote the current occupancy map, constructed from onboard sensor measurements $\mathbb{S}$ and odometry $\mathbb{O}$. The map consists of voxels $m\in\mathbb{M}$, where each voxel is categorized as free ($m\in\mathbb{M}_{free}$), occupied ($m\in\mathbb{M}_{occupied}$), or unknown ($m\in\mathbb{M}_{unknown}$). During robotic exploration we seek to recover the complete map $\mathbb{M}_{comp}$ of a target region $\mathbb{M}_{target}$ from successive measurements from the onboard sensors. While accurate completion and alignment of $\mathbb{M}_{comp}$ with $\mathbb{M}_{target}$ is the primary goal, secondary goals include minimizing exploration time and maximizing volumetric gain per second, provided safety is not compromised.
@@ 3.2 Frontier Identification and Evaluation
We identify and evaluate frontiers of interest in $\mathbb{M}$ that can enhance the robot’s understanding of the environment during exploration. In general, “interesting” frontiers will maximize the number of unknown voxels available for occupancy prediction while considering common exploration metrics such as directionality, distance from target, and reachability [biggie2023flexible, dang2020graph]. These frontiers guide the robot’s exploration.
@@ 3.3 Dense Occupancy Prediction
Dense occupancy prediction estimates the probability of occupancy for each voxel $m\in\mathbb{M}$, producing a continuous value in $[0,1]$ where 0 indicates free and 1 indicates occupied. These predictions are used to hypothesize structure beyond the currently observed map generated by the robots sensor measurements.
@@ 3.4 Forward Diffusion
Let $x_{0}$ denote an occupancy subgrid ($x\subset\mathbb{M}$), sampled from the data distribution $q(x_{0})$. By sampling from the data distribution $x_{0}\sim q(x_{0})$, the forward diffusion process is defined as a Markov chain of variables $x_{1},...,x_{T}$ that iteratively adds Gaussian noise to the sample. A diffusion step at time $t$ in this chain is defined as:
@@
Conveniently, we can apply the reparameterization trick to directly sample $x_{t}$ given $x_{0}$ using the conditional distribution:
@@ 3.5 Reverse Diffusion
Reverse diffusion defines a generative process that inverts the forward noising procedure. Specifically, it is modeled as a Markov chain of learned Gaussian transitions, parameterized by a neural network with learnable parameters $\theta$:
@@
where $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the predicted mean and covariance of the Gaussian distribution over $x_{t-1}$. The parameters $\theta$ correspond to the SceneSense diffusion model, trained to recover occupancy maps from noisy inputs.
Given the initial state of a noisy occupancy map from a standard multivariate Gaussian distribution $x_{t}\sim\mathcal{N}(0,I)$, the reverse diffusion process iteratively predicts $x_{t-1}$ at each time step $t$ until reaching the final state $x_{0}$ which is the goal occupancy map. Similar to the Markov chain defined forward diffusion process the joint distribution of the reverse diffusion process is simply the product of the applied learned Gaussian transitions $p_{\theta}(x_{t-1}|x_{t})$: